From: Jet Propulsion Laboratory
Posted: Wednesday, October 3, 2001
Wednesday, October 3, 2001 (DOY 269/19:00:00 to DOY 276/19:00:00 UTC)
Launch / Days since Launch = Nov. 7, 1996 / 1792 days
Start of Mapping / Days since Start of Mapping = April 1, 1999 / 916 days
Total Mapping Orbits = 11,498
Total Orbits = 13,181
Background Sequences - The spacecraft is operating nominally in performing daily recording and transmission of science data. The mm153 sequence has performed well since it started 01-256 (9/13/01). It terminates on 01-290 (10/17/01).
Other - Three Roll Only Targeted Observations (ROTOs) were performed this past week. All three occurred during scheduled communications orbits. Previously, MGS had been restricted to performing ROTOs only during non-communications orbits. This new capability of performing ROTOs during communications orbits will allow MGS to perform ROTOs during the Odyssey aerobraking support phase, when MGS has 24 hours of DSN coverage each day. MGS has completed 201 ROTOs to date.
All subsystems report good health and status.
The MOLA instrument team has completed their investigation of the MOLA anomaly of 01-181 (6/30/01). The collective results from the instrument tests indicate the anomaly was caused by an interruption of the 100-MHz signal from the precision oscillator in the MOLA altimeter electronics. Without the oscillator signal, the 10-Hz laser trigger signal cannot occur and the laser cannot fire. All testing since July, 2001 shows that the instrument responds properly to commanding and that all other parts of the MOLA are operating normally. The diagnostic tests also demonstrated that the MOLA is fully capable of continuing to collect data in passive radiometer mode.
There have been 19 uplinks to the spacecraft during the past week, including new star catalogs and ephemeris files, instrument command loads, and the MZ137 and MZ138 ROTO mini-sequences. 5,790 command files have been radiated to the spacecraft since launch.
A Solar Array (SA) and High Gain Antenna (HGA) flexible modes baseline diagnostic test will be performed on 01-278 (10/05/01). It is designated MZ139. Performing this test periodically should give us insight into any additional SA hinge or HGA gimbal degradation, if it occurs.
Five MOC defocus calibrations will be performed between 01-281 (10/8/01) and 01-289 (10/16/01).
A fifth and final MOLA diagnostic test will be conducted on 01-283 (10/10/01) after which the MOLA will be left on to operate in passive radiometer mode.
// end //